# Moto: Latent Motion Token as the Bridging Language for Robot Manipulation

### v1

@@ Abstract
footnotetext: Corresponding Authors.
@@ 1 Introduction
In this work, we introduce Moto, which utilizes Latent Motion Tokens as a bridging “language” to model visual motions between video frames in an unsupervised manner. As illustrated in Fig. 1, we first train a discrete Latent Motion Tokenizer to produce compact latent motion tokens that capture dynamics between video frames without external supervision. We then pre-train Moto-GPT using a GPT-based architecture to predict the next latent motion token, absorbing motion priors from videos. These learned priors are subsequently transferred to enhance robot manipulation tasks through a co-fine-tuning strategy.
@@
We conduct extensive experiments to validate our claims from various perspectives: (1) Latent Motion Token as an Interpretable Motion Language: Experiments show that latent motion tokens encapsulate compact and expressive representations of motion, effectively reconstructing and understanding motion trajectories in videos. (2) Pre-trained Moto-GPT as a Useful Motion Prior Learner: Results indicate that the pre-trained Moto-GPT achieves promising outcomes in predicting plausible motion trajectories and assessing the rationality of robot trajectories based on output likelihood. (3) Fine-tuned Moto-GPT as an Effective Robot Policy: The fine-tuned Moto-GPT demonstrates significant performance improvements over counterparts trained without motion priors, especially with limited training data, highlighting its effectiveness in transferring learned motion knowledge to robot manipulations.
In summary, our contributions are threefold as below:
@@
We believe the vast reservoir of interaction-rich knowledge in video data presents a crucial opportunity for advancing robot learning and hope this paper inspires further exploration of effective autoregressive representations for acquiring valuable priors through pre-training, ultimately enhancing robotic capabilities.
@@ 2 Related Work
Vision-Language-Action Models.
Recent studies have increasingly employed transformers as unified vision-language-action (VLA) architectures to generate robot actions from sequential observations and language instructions [5, 48, 25]. Inspired by the success of pre-training in vision-language transformers [6, 44, 1, 36], VLA model pre-training has gained traction. One approach fine-tunes policy models from powerful vision-language models pre-trained on large image-text datasets [16, 62, 32]. Another explores training generalist policy models on diverse cross-embodiment robot data with action labels [42, 28, 15, 52]. In contrast, our work aims to enhance VLA models through generative pre-training on video data, which offers richer interaction details than text and images and requires no hardware-specific labels of low-level robot actions. Beyond VLA models, several contributions focus on improving robot manipulation performance. Some extend input observations from single-view RGB images to include multi-perspective views and depth information [35, 8, 59]. Techniques like action chunking and policy diffusion also enhance action precision [22, 13, 27]. Additionally, some works [20, 34] decompose high-level language instructions into latent skills learned through auxiliary training objectives during imitation learning.
Robot Learning from Videos
Videos provide rich knowledge about physical dynamics, making them ideal for robot learning. Early works [43, 38] utilized contrastive learning with egocentric videos to enhance visual representations for manipulation. Some studies [17, 29, 33, 4, 30] generate videos or images as intermediate plans for guiding low-level control. Recent research [54, 9, 23] has shifted towards generative video pre-training followed by fine-tuning to create end-to-end policy models. Escontrela et al. [19] pre-trains an autoregressive video prediction model to provide reward signals for reinforcement learning. These works primarily use pixel values or patch-level tokens of video frames as their pretraining target. In contrast, our approach focuses on latent motion tokens as prediction targets, emphasizing key visual motions while decoupling irrelevant details. Additionally, some studies build world models through action-conditioned video generation [21, 56, 55], facilitating reinforcement learning or serving as interactive environments. Notably, Genie [7] proposes unsupervised learning of latent actions from large-scale videos to create a versatile 2D gaming simulator. Our goal, however, is to train a generalized policy model for robot manipulation, which is more complex than developing a 2D gaming simulation environment. Concurrently, Ye et al. [58] pre-train a policy model to predict one-step future latent actions, while Chen et al. [12] use latent actions as intermediate goals for low-level policies. Our approach differs by pre-training an end-to-end policy model to autoregressively predict a trajectory of latent motion tokens for future video clips.
@@ 3.1 Overview
Moto utilizes autoregressive generative pre-training on latent motion token sequences to learn motion priors from videos, followed by co-fine-tuning on action-labeled data for robot control. As illustrated in Figure 2, Moto consists of three stages: 1) unsupervised training of the Latent Motion Tokenizer, 2) pre-training of the generative model Moto-GPT, and 3) co-fine-tuning for robot action policy. In Sec 3.2, we detail the Latent Motion Tokenizer, which encodes visual dynamics into quantized latent motion tokens. We also describe the training procedures for Moto-GPT, including motion token autoregressive pre-training in Sec 3.3 and supervised co-fine-tuning in Sec 3.4. Implementation details can be found in the Supplementary Material.
@@ 3.2 Latent Motion Tokenizer
The Latent Motion Tokenizer, as shown in Figure 3, learns a latent “language” to capture essential visual motions between successive video frames [^1] To ensure significant visual differences, we down-sample the original video by a certain rate. in an unsupervised manner. The architecture follows a standard auto-encoder design for motion tokenization and detokenization. The tokenization employs an M-Former, a multi-layer transformer that extracts motion features from the last-layer patch features of the current frame $o_{t}$ and the preceding frame $o_{t-1}$ using a frozen pre-trained ViT encoder [24]. We concatenate 8 learnable query embeddings with these patch features as additional input to the M-Former, where the queries interact through self-attention layers. The output query features are then processed by a VQ codebook with a vocabulary size of 128 to produce discrete latent motion tokens.
@@
Figure 3: The Latent Motion Tokenizer produces discrete motion tokens from two consecutive video frames. It regularizes the decoder to reconstruct the second frame based on the first one and the discrete tokens, effectively capturing the motion between frames.
@@ 3.3 Motion Token Autoregressive Pre-training
With the Latent Motion Tokenizer, Moto-GPT is allowed to learn about diverse visual motions from videos, using latent motion tokens as a bridging language. As shown in Figure 2, Moto-GPT is pre-trained with a next-motion-token prediction objective. For a video clip $[o_{0},o_{1},...,o_{T}]$, we derive a chunk of latent motion tokens for each pair of consecutive frames, concatenating them chronologically to form a sequence. Moto-GPT employs a GPT-style transformer for autoregression on these motion token trajectories. Additionally, we prepend the text features from the instruction and the visual features from the initial video frame as input prompts. The pre-training objective maximizes the likelihood of the ground-truth latent motion token sequence given the language instruction and the initial video frame:
@@
where $\boldsymbol{l}$ and $\boldsymbol{v}$ are text and visual features from the frozen pre-trained T5 [47] and ViT [24] models, respectively. $\boldsymbol{m}_{<i}$ represents the latent motion tokens preceding the current token $m_{i}$, and $\boldsymbol{\Theta}$ denotes the trainable model parameters. Here, $M=K*T$, where $K$ is the number of tokens for motion between successive frames and $T$ is the video length.
@@ 3.4 Co-fine-tuning for Robot Manipulation
To address this, during fine-tuning, we introduce special action query tokens into Moto-GPT’s input, enabling the generation of real robot actions through a flexible action head, as illustrated in the right part of Figure 2. Specifically, $N$ query tokens are added after the latent motion token chunk at each time step, where $N$ corresponds to the number of robot actions occurring between two video frames. The fine-tuning stage follows the same causal mask mechanism as pre-training in general. Nevertheless, the latent motion tokens do not attend to the newly inserted action query tokens to stay consistent with the pre-training setting. Besides, we randomly mask 50% of the attention from action query tokens to latent motion tokens, allowing knowledge transfer while reducing dependency on ground-truth conditions. This also improves inference efficiency, enabling direct queries to Moto for real actions without generating latent motion tokens. This can be achieved by using padding tokens as placeholders for latent action tokens, blocking attention from action query tokens to these placeholders.
@@
4 Experiment Setup
4.1 Benchmarks and Datasets
We use SIMPLER [31] and CALVIN [40] as the main evaluation benchmarks for robot manipulation.

### v4

# Moto: Latent Motion Token as the Bridging Language for Learning Robot Manipulation from Videos

@@ Abstract
**footnotetext: Part of the work done during internship at Tencent ARC Lab.footnotetext: Corresponding Authors.
@@ 1 Introduction
In this work, we introduce Moto, which utilizes Latent Motion Tokens as a bridging “language” to model visual motions between video frames in an unsupervised manner. As illustrated in Fig. 1, we first train a discrete Latent Motion Tokenizer to produce compact latent motion tokens that capture dynamics between video frames without external supervision. We then pre-train Moto-GPT using a GPT-based architecture to predict the next latent motion token, absorbing motion priors from videos. These learned priors are subsequently transferred to enhance robot manipulation performance through a co-fine-tuning strategy.
@@
We conduct extensive experiments to validate our claims from various perspectives: (1) Latent Motion Tokens as an Interpretable Motion Language: Experiments show that latent motion tokens encapsulate compact and expressive representations of visual motions and exhibit promising cross-embodiment transfer ability (even from human to robot). (2) Pre-trained Moto-GPT as a Useful Motion Prior Learner: Results indicate that pre-trained Moto-GPT achieves promising outcomes in predicting plausible motion trajectories and assessing the rationality of robot trajectories based on output likelihood. (3) Fine-tuned Moto-GPT as an Effective Robot Policy: The fine-tuned Moto-GPT demonstrates significant performance gains over counterparts trained from scratch without motion priors, especially with limited training data. The performance can be further boosted with human video pre-training, highlighting the potential of our approach in transferring motion knowledge learned from Internet-scale videos to robot manipulation.
In summary, our contributions are as below:
@@
  • We conduct systematic experiments and analyses to validate each training stage of our approach and the effectiveness of the resulting robot policy. Our open-source code provides clear guidance for future exploration.
@@ 2 Related Work
Vision-Language-Action Models. Recent studies have utilized transformers as vision-language-action (VLA) architectures to generate robot actions from observations and language instructions [4, 47, 25, 30]. VLA model pre-training has gained traction, with approaches either fine-tuning policy models from powerful vision-language models pre-trained on large image-text datasets [15, 62, 32] or training generalist models on diverse robot data with action labels [41, 27, 14, 52, 2]. Our work enhances VLA models through generative pre-training on video data, providing richer interaction details. Beyond VLA models, several studies improve robot manipulation by incorporating multi-perspective views and depth information [35, 7, 59], and employing techniques like action chunking and policy diffusion for precision [22, 11, 26]. Additionally, some works [19, 34] decompose high-level language instructions into latent skills via auxiliary training objectives.
Robot Learning from Videos. Early works [42, 37] used contrastive learning with egocentric videos to enhance visual representations. Some studies [16, 28, 33, 3, 29] generate videos as intermediate plans for low-level control. Recent research [54, 8, 23] has focused on generative video pre-training for end-to-end policy models. Escontrela et al. [18] pre-train autoregressive video prediction models for reinforcement learning. These works primarily use pixel values or patch-level tokens. In contrast, our approach targets latent motion tokens, emphasizing key visual motions. Some studies build world models through action-conditioned video generation [21, 56, 55], aiding reinforcement learning. Genie [6] introduces unsupervised learning of latent actions from large-scale videos for 2D gaming simulators. Similar latent action concepts are explored in LAPO [49] and DynaMo [12], though LAPO is limited to video games and DynaMo focuses on visual representation learning. Concurrently, LAPA [58] is pre-trained to predict one-step future latent action, while IGOR [10] uses latent actions as intermediate goals. Our work differs by pre-training an end-to-end policy model to autoregressively predict a trajectory of latent motion tokens for future video clips, providing a more natural modeling of sequential visual motions from videos. Additionally, we conduct a rigorous analyses to validate the effectiveness of policy and each training stage, offering clear guidance for future research.
@@ 3.1 Overview
Moto utilizes autoregressive generative pre-training on latent motion token sequences to learn motion priors from videos, followed by co-fine-tuning on action-labeled data for robot control. As illustrated in Fig. 2, Moto consists of three stages: 1) unsupervised training of the Latent Motion Tokenizer, 2) pre-training of the generative model Moto-GPT, and 3) co-fine-tuning for robot action policy. In Sec. 3.2, we detail the Latent Motion Tokenizer, which encodes visual dynamics into quantized latent motion tokens. We also describe the training procedures for Moto-GPT, including motion token autoregressive pre-training in Sec. 3.3 and supervised co-fine-tuning in Sec. 3.4. Implementation details can be found in the Supplementary Material.
@@ 3.2 Latent Motion Tokenizer
The Latent Motion Tokenizer, as shown in Fig. 2, learns a latent “language” to capture essential visual motions between successive video frames [^1] To ensure significant visual differences, we down-sample the original video by a certain rate. in an unsupervised manner. The architecture follows a standard auto-encoder design for motion tokenization and detokenization. The tokenization employs an M-Former, a multi-layer transformer that extracts motion features from the last-layer patch features of the current frame $o_{t}$ and the preceding frame $o_{t-1}$ using a frozen pre-trained ViT encoder [24]. We concatenate 8 learnable query embeddings with these patch features as additional input to the M-Former, where the queries interact through self-attention layers. The output query features are then processed by a VQ codebook with a vocabulary size of 128 to produce discrete latent motion tokens.
@@ 3.3 Motion Token Autoregressive Pre-training
With the Latent Motion Tokenizer, Moto-GPT is allowed to learn about diverse visual motions from videos, using latent motion tokens as a bridging language. As shown in Fig. 2, Moto-GPT is pre-trained with a next-motion-token prediction objective. For a video clip $o_{0},o_{1},...,o_{T}]$, we derive a chunk of latent motion tokens for each pair of consecutive frames, concatenating them chronologically to form a sequence. Moto-GPT employs a GPT-style transformer for autoregression on these motion token trajectories. Additionally, we prepend the text features from the instruction and the visual features from the initial video frame as input prompts. The pre-training objective maximizes the likelihood of the ground-truth latent motion token sequence given the language instruction and the initial video frame:
@@
where $\boldsymbol{l}$ and $\boldsymbol{v}$ are text and visual features from the frozen pre-trained T5 [46] and ViT [24] models, respectively. $\boldsymbol{m}_{<i}$ represents the latent motion tokens preceding the current token $m_{i}$, and $\boldsymbol{\Theta}$ denotes the trainable model parameters. Here, M=KTM=K*T, where $K$ is the number of tokens for motion between successive frames and $T$ is the video length.
@@ 3.4 Co-fine-tuning for Robot Manipulation
To address this, during fine-tuning, we introduce special action query tokens into Moto-GPT’s input, enabling the generation of real robot actions through a flexible action head, as illustrated in the right part of Fig. 2. Specifically, $N$ query tokens are added after the latent motion token chunk at each time step, where $N$ corresponds to the number of robot actions occurring between two video frames. The fine-tuning stage follows the same causal mask mechanism as pre-training in general. Nevertheless, the latent motion tokens do not attend to the newly inserted action query tokens to stay consistent with the pre-training setting. Besides, we randomly mask 50% of the attention from action query tokens to latent motion tokens, allowing knowledge transfer while reducing dependency on ground-truth conditions. This also improves inference efficiency, enabling direct queries to Moto for real actions without generating latent motion tokens. This can be achieved by using padding tokens as placeholders for latent action tokens, blocking attention from action query tokens to these placeholders.
Figure 3: Illustration of real-world evaluation tasks.
@@
Figure 4: Interpretability of latent motion tokens. Each row displays reconstructed frames from the same initial frame using different latent motion tokens, while each column shows frames reconstructed from the same latent motion tokens with varying initial frames. The latent motion tokens exhibit consistent (see columns) and discriminative (see rows) semantics, despite being trained in an unsupervised manner.
Figure 5: Video imitation generation via latent motion tokens, where a sequence of motion tokens extracted from a demonstration video are decoded into a new video. This generated video is based on a different initial frame while preserving the original movement semantics.
4 Benchmarks and Datasets
More details of experimental settings can be found in the supplementary material.
SIMPLER [31]. We focus on three tasks with the Google Everyday Robot: “pick coke can”, “move near”, and “open/close drawer”. For pre-training, we leverage a subset of Open-X-Embodiment (OXE) [52], consisting of 109k trajectory videos [4, 53, 48, 40, 13, 36, 38, 44, 61, 9, 60]. Fine-tuning is performed using 73k action-labeled expert trajectories from the RT-1 Robot-Action dataset [4].
[…]
Real-world Robot Experiments. We conduct real-world evaluations with a FANUC LR Mate 200iD robot on three tasks: “pick-place banana”, “close laptop”, and “disassembly” (Fig. 3). Pre-training utilizes OXE data, and we collect 90 teleoperated demonstrations (30 per task) for fine-tuning. Each task is tested 10 times with randomized object positions. Generalizability is evaluated in two scenarios: (i) Novel Object, altering the object’s color, texture, and shape; (ii) Visual Distractor, introducing irrelevant objects.
5 Experiments
5.1 Latent Motion Token as an Interpretable Motion Language
The Latent Motion Tokenizer effectively reconstructs authentic next frames using ground-truth latent motion tokens, capturing key dynamics between initial and subsequent frames, as evidenced by supplementary examples. This demonstrates its ability to represent fine-grained motion details, with its decoder acting as a reliable simulator for visualizing environmental changes. Fig. 4 highlights the controllability and consistency of latent motion tokens. Different token chunks produce varied motion orientations and scales, while identical chunks yield consistent effects across different starting observations with varied robot embodiments. By concatenating token chunks from consecutive frames, motion trajectories can be represented sequentially, akin to natural language, enabling contextualized motion generation across diverse initial observations (Fig. 5). This underscores the potential of latent motion tokens as a unified language for imitation learning.
Quantitatively, Table 1 shows the semantic interpretability of latent motion tokens. A video classifier using initial frame ViT patch features and concatenated latent motion tokens for seven subsequent frames achieves 79.7% accuracy in predicting semantic labels for 34 CALVIN tasks. This performance is comparable to using ViT features for all eight frames, despite reducing input features from 196 to 8 tokens per frame, confirming that latent motion tokens provide a compact, expressive, and interpretable representation of visual motions linked to high-level semantics.
5.2 Moto-GPT as a Useful Motion Prior Learner
Video Representation Semantic Acc.
Initial frame 0.292
Initial frame repeated by 8 times 0.283
Initial frame + 7 subsequent frames 0.828
Initial frame + 7 latent motion token chunks 0.797
Moto-GPT is pre-trained through autoregression on videos using latent motion tokens, enabling it to predict motion trajectories from initial observations and diverse language prompts, as shown in Fig. 6. The top-k token prediction accuracy and the visualization of predicted video trajectories for more complex actions (e.g., “rotation” and “stacking”) are provided in the supplementary material. This demonstrates Moto-GPT’s ability to acquire prior knowledge essential for robot action inference based on human instructions. Additionally, latent motion tokens allow Moto-GPT to interpret trajectory videos as compact token sequences and evaluate their rationality through the autoregressive likelihood (Eq. 1). Fig. 7 illustrates the potential of using Moto’s log-likelihoods as a reward signal for trajectory videos, indicating how well a trajectory aligns with Moto-GPT’s distribution and measuring the temporal consistency of behavior. To assess this, we collected 98 video triplets in CALVIN using the baseline policies and a random policy. Each triplet consists of three types of trajectory videos originating from the same environment state. The averaged log-likelihoods for each trajectory type at each sequence step, shown in Fig. 7, clearly differentiate successful trajectories from failures and random attempts.
5.3 Moto-GPT as an Effective Robot Policy
Method Pick Coke Can Move Near Open / Close Drawer Overall
RT-1-X [4] 0.820 0.330 0.550 0.567 0.317 0.296 0.891 0.597 0.534
RT-2-X [62] 0.740 0.740 0.880 0.787 0.779 0.157 0.343 0.250 0.607
Octo-Base [41] 0.210 0.210 0.090 0.170 0.042 0.009 0.444 0.227 0.169
OpenVLA [27] 0.270 0.030 0.190 0.163 0.462 0.194 0.518 0.356 0.248
OpenVLA (fine-tuned) [27] 0.470 0.080 0.540 0.363 0.542 0.102 0.361 0.231 0.349
Moto 0.820 0.500 0.900 0.740 0.604 0.130 0.732 0.431 0.614
[…]
Model Observation Space Tasks competed in a row (1000 chains)
1 2 3 4 5 Avg. Len.
SuSIE [3] Static RGB 0.870 0.690 0.490 0.380 0.260 2.69
Moto Static RGB 0.897 0.729 0.601 0.484 0.386 3.10
Moto w/o Motion Token Static RGB 0.779 0.555 0.380 0.256 0.167 2.14
Overall Performance. After fine-tuning, Moto-GPT [^3] Moto-GPT is referred to as Moto in the following tables and figures. was evaluated on the SIMPLER and CALVIN benchmarks, demonstrating promising results as shown in Tables 2 and 3. It significantly outperforms Moto w/o Motion Token, which is trained from scratch without latent motion tokens, underscoring the effectiveness of transferring motion priors learned from videos to enhance robot manipulation tasks. Despite having only 98M parameters for the GPT backbone and no access to action labels during pre-training, Moto-GPT performs comparably to larger models like RT-2-X (PaLI-X 55B) and OpenVLA (Prismatic-7B) on SIMPLER. It also maintains competitiveness against OpenVLA (fine-tuned), which is further fine-tuned specially on the RT-1 Robot-Action trajectories, despite its pre-training data already containing action labels from this dataset. Moto-GPT also generalizes well in the unseen CALVIN environment, outperforming baseline models that use various pre-training strategies (see supplementary material for detailed descriptions about baselines). Notably, Moto-GPT relies solely on RGB images from a static camera, achieving competitive results compared to GR-1, which is pre-trained to predict future pixel values and uses gripper RGB and proprioceptive states as additional inputs. Our findings suggest that focusing on motion dynamics rather than frame-level details is a more effective approach for learning from videos.
Figure 8: Evaluation results in the real-world environment.
Real-world Experiments. We additionally test Moto-GPT on three real-world tasks. As shown in Fig. 8, Moto-GPT consistently outperforms Moto w/o Motion Token on these tasks, improving the average success rate from 23.33% to 60%. For generalizability evaluation, we add visual distractors (Visual Distractor) or change the appearance of manipulated objects (Novel Object). Moto-GPT still enhances the average performance by 20% and 30% under varying conditions. This further demonstrates the robustness of Moto-GPT in real-world deployment.
Figure 9: With additional human video (SSV2) pre-training, Moto (OXE+SSV2) significantly outperforms both Moto w/o Motion Token and Moto (OXE) on the Move Near task in SIMPLER.
Figure 10: Visualization of motion transfer from human to robot videos via latent motion tokens.
Learning from Human Videos. We expand our approach to Internet-scale human videos, demonstrating their potential for large-scale robotics pre-training. Using SSV2 [20] capturing simple human hand movements, we filter out less motion-relevant videos, retaining 105k human videos for a preliminary study. These are combined with 109k OXE robot videos to train the Latent Motion Tokenizer and pre-train Moto-GPT. Fig. 9 shows that adding human videos further enhances Moto-GPT’s performance compared with Moto (OXE) which is only pre-trained on OXE videos. Additionally, latent motion tokens can act as a unified “language”s to translate human motions into semantically aligned robot actions (Fig. 10). This enables not only pre-training with human videos but also in-context robot learning guided by online human demonstrations. Future work will improve model architectures and incorporate more diverse human videos to tackle complex manipulation tasks.
Data Efficiency. Moto-GPT is pre-trained exclusively on videos, bypassing the need for supervised data with action labels. This enables pre-training on large, easily accessible video datasets, followed by fine-tuning with smaller action-labeled trajectories for policy adaptation. Fig. 11 shows that Moto-GPT fine-tuned with varying amounts of labeled data consistently outperforms its variant trained from scratch without latent motion tokens, especially with limited data. For instance, Moto-GPT achieves a 52.5% success rate with just 1% of labeled data, compared to 0% for the variant. This demonstrates Moto-GPT’s efficiency in action adaptation and its potential to improve robot manipulation tasks through large-scale video pre-training.
[…]
Ablations on Policy Fine-tuning Methods. In Fig. 12, we assess Moto’s co-fine-tuning strategy. Moto-IML and Moto-DM use the same pre-training as Moto-GPT but differ in fine-tuning: Moto-IML excludes the loss term for latent motion token prediction, while Moto-DM omits latent motion tokens from the input entirely. Compared to Moto w/o Motion Tokens, which is trained from scratch, both Moto-IML and Moto-DM show improved performance due to pre-training motion priors, yet they still lag behind Moto-GPT. This highlights the importance of retaining latent motion tokens in the sequence, allowing action query tokens to transfer knowledge through direct attention. Furthermore, co-fine-tuning for latent motion token prediction helps preserve the learned motion priors in Moto-GPT.
[…]
6 Conclusion and Discussion
We present Moto, a method that utilizes latent motion tokens as a “language” interface to bridge generative pre-training on video data with precise robot control. By learning motion-related priors from videos without the need for hardware-specific action labels, Moto effectively translates learned motions into actual robot actions. We demonstrate that Moto can not only be used for pre-training on robot videos, but also be extended to Internet-scale human videos and enable cross-embodiment knowledge transfer. Future research should aim to scale up the pre-training dataset and optimize fine-tuning to enhance performance across a broader range of robotic applications.
Acknowledgement
The research work described in this paper was conducted in the JC STEM Lab of Autonomous Intelligent Systems funded by The Hong Kong Jockey Club Charities Trust.
Supplementary Material
7 Details on Experiment Setup
7.1 Benchmarks